OpenStudy (anonymous):

how to solve 3 ln 5 + 5lnx

OpenStudy (callisto):

I think you need an equal sign to 'solve' ?

OpenStudy (anonymous):

maybe the question is "write as a single log" which is not solving for x, although i guess it is "solve the problem"

OpenStudy (anonymous):

what does the question ask?

OpenStudy (anonymous):

it syas write the expression as a single natural logarithm

OpenStudy (anonymous):

ah in that case take two steps. first use the rule \[n\log(a)=\log(a^n)\] to write both terms, and then use \[\log(a)+\log(b)=\log(ab)\] to write as a simgle logarithm